The playing field for a particular sport is a rectangle whose length is 5 feet more than twice the width. The perimeter of the playing field is 196 feet. Find the dimensions of the playing field.
Found 3 solutions by MowMow, ikleyn, greenestamps:
Answer by MowMow(42)

Let w be the width of the field,
Then w+5 is the length of the field.
The perimeter = 196 = 2w + 2(w+5)
p = 186 = 4w
The width = 4.5 ft. and the length = 9.5 ft.

Answer by ikleyn(52794)

.

            What you see below is my correct solution to the problem,

            produced against the incorrect solution by @MowMow.


Let w be the width of the field,

Then 2w+5 is the length of the field.

The perimeter = 196 = 2w + 2(2w+5)

186 = 6w

The width = 186/6 = 31 ft. and the length = 2*31+5 = 67 ft.     ANSWER

-----------

Solved and completed,

to disproof the incorrect solution by @MowMow.



Answer by greenestamps(13200)



For tutor @ikleyn....

The length is 5 more than TWICE the width....

----------------------------
Oh, now I see you have already corrected your response....